# Unified Source-Free Domain Adaptation

### v1

@@ Abstract
In the pursuit of transferring a source model to a target domain without access to the source training data, Source-Free Domain Adaptation (SFDA) has been extensively explored across various scenarios, including closed-set, open-set, partial-set, and generalized settings. Existing methods, focusing on specific scenarios, not only address only a subset of challenges but also necessitate prior knowledge of the target domain, significantly limiting their practical utility and deployability. In light of these considerations, we introduce a more practical yet challenging problem, termed unified SFDA, which comprehensively incorporates all specific scenarios in a unified manner. To tackle this unified SFDA problem, we propose a novel approach called Latent Causal Factors Discovery (LCFD). In contrast to previous alternatives that emphasize learning the statistical description of reality, we formulate LCFD from a causality perspective. The objective is to uncover the causal relationships between latent variables and model decisions, enhancing the reliability and robustness of the learned model against domain shifts. To integrate extensive world knowledge, we leverage a pre-trained vision-language model such as CLIP. This aids in the formation and discovery of latent causal factors in the absence of supervision in the variation of distribution and semantics, coupled with a newly designed information bottleneck with theoretical guarantees. Extensive experiments demonstrate that LCFD can achieve new state-of-the-art results in distinct SFDA settings, as well as source-free out-of-distribution generalization. Our code and data are available at https://github.com/tntek/source-free-domain-adaptation.
@@ 1 Introduction
Due to escalating demands for privacy and information protection in public security and commercial competition, well-annotated training data are recognized as crucial for varying purposes, requiring stringent access control. However, in practical scenarios, the use of a source model (pre-trained on the source domain) without access to the actual source data, referred to as Source-Free Domain Adaptation (SFDA) [1], has become more feasible. This approach involves the challenges posed by strict data access controls and emphasizes the adaptation of a pre-trained model to new domains without relying on the availability of the original source data.
SFDA has been explored across various scenarios, as summarized in Table I. The most constrained scenario, termed closed-set, assumes identical categories of interest in both the source and target domains, focusing on addressing covariate shift, such as domain-specific appearance distributions [2]. Generalized SFDA extends this by requiring the adapted model to perform well in both the source and target domains without forgetting the source domain [3]. Moving beyond the vanilla category identity constraint, open-set SFDA [4] considers the target domain with additional categories, while partial-set SFDA [5] takes the opposite direction. Both scenarios require handling the additional challenge of semantic shift.
@@
TABLE I: Summary of different SFDA settings. $C_{s}$/$C_{t}$: The category set of the source/target domain.
@@
To address the aforementioned limitation, this study introduces a more realistic a more realistic yet challenging problem, referred to as Unified Source-Free Domain Adaptation (Unified SFDA), with the goal of comprehensively addressing all the specific scenarios mentioned earlier in a unified manner. To achieve this, we propose a novel approach called Latent Causal Factors Discovery (LCFD), going beyond the conventional statistical association learning of related variables by exploring the underlying causal relationships [10]. The discovered causal mechanism is expected to be more reliable under varying distributions and semantic contexts, providing a unified solution for distinct SFDA scenarios.
In the absence of label supervision for both distributional and semantic variations, our LCFD is specifically formulated using a structural causal model in the logits space. The latent causal factors are disentangled into two complementary parts: (i) external causal factors and (ii) internal causal factors, consider that the information of both domain is not necessarily complete. For the former, we leverage a pre-trained large Vision-Language (ViL) model with rich knowledge such as CLIP [11], which has been exposed to a vast amount of multimodality information sources. The latter is identified under the guidance of discovered external causal factors. The discovery of external and internal causal factors is alternated by designing a self-supervised information bottleneck with theoretical guarantees.
@@
    We introduce a novel approach for unified SFDA, termed LCFD. Instead of learning the statistical description of the problem reality as conventional methods do, our LCFD is formulated under the causality perspective. It aims to reveal the underlying causal relationships between latent variables and model decisions, providing favorable robustness against both distributional and semantic shifts.
@@
    Extensive experiments demonstrate that LCFD consistently achieves new state-of-the-art results on varying SFDA scenarios as well as source-free out-of-distribution generalization.
@@ 2.1 Source-Free Domain Adaptation
Most of the prior SFDA methods consider the closed-set setting where the source and target domains share the same classes and the focus is cross-domain distribution alignment. Existing methods exist in two categories. The first converts SFDA to Unsupervised Domain Adaptation (UDA) by constructing a pseudo-source domain [12], exploiting source prototype-guided data generation [13] or splitting source-like subset from target data [14]. The second follows the paradigm of self-supervised learning [15, 16], introducing self-guidance. Besides widely used pseudo-labels [15, 17], geometry information [18, 19, 20], constractive data [9] and historical data [8] have also been exploited.
Recently, more practical settings have been studied. For example, the generalized setting shifts aims to mitigate forgetting the source domain [3]. A representative approach is combining continue learning techniques with cross-domain adaptation, e.g., the domain attention-based gradient regulation [3] and source guidance constructed by historical information refining [8]. There exist works considering open-set [7] and partial-set [21] settings. For instance, to control the open-set risk, a progressive balanced pseudo-labeling strategy [7] and closed-set class prototypes are proposed [22]. Estimating the target data distribution helps with negative transfer reduction from the partial classes shift [23].
@@
Existing works as above focus on a specific SFDA setting, limiting their usability and generality substantially in practice. To address this limitation, we introduce a unified SFDA setting that incorporates all the pre-existing scenarios. With this new setup, we aim to foster more advanced SFDA methods that could generalize to a variety of problem settings without the need for designing and maintaining an array of distinct algorithms.
@@ 2.2 Causality methods in transfer learning
Causality methods aim for a robust function relation between random variables, which is insensitive to the extra variation/disturbing [27]. This property leads to recent popularity of this new paradigm in transfer learning, e.g., UDA [28], domain generalization (DG) [29] and out of distribution (OOD) [30]. These attempts aim to reduce non-causal factors by introducing artificial intervention. For example, the probability distribution change (used to represent the domain shift) and semantic consistency constraint are jointly employed to implement this intervention. As source labels are available to ensure consistency, these methods put the highlight on constructing the distribution change, following in two lines. One explicitly generates augmented data to disturb the distribution, e.g., the non-linear augmentation and spatially-variable blending augmentation [31], inverse Fourier transformation augmentation [32], and generative models-based cross-domain image style transformation [33, 34]. The other implicitly modifies the distribution by exploiting cross-domain data. For example, MatchDG [35] developed cross-domain contrastive learning where input’s positive sample was randomly selected from samples with the same class. COR [30] adopted the variational inference encoder to infer unobserved causal factors from historical data, taking the time domain as a natural distribution change.
Compared with the existing works, the following three features distinguish LCFD from them. (i) LCFD does not rely on real labels that are indispensable for previous works. (ii) LCFD discovers the causal factors instead of indirectly removing the non-causal ones. (iii) LCFD builds the structural causal model at the logits level whilst previous methods construct on raw data.
@@ 3.1 Unified SFDA problem
SFDA aims to transfer a model pre-trained on the source domain to a different but related target domain without labelling. Formally, let $\mathcal{X}_{s}\!=\!\{\boldsymbol{x}_{i}^{s}\}_{i=1}^{n_{s}}$ and $\mathcal{Y}_{s}\!=\!\{{y}_{i}^{s}\}_{i=1}^{n_{s}}$ be the source samples and their truth labels. Similarly, the unlabeled samples and the truth target labels are $\mathcal{X}\!=\!\{\boldsymbol{x}_{i}\}_{i=1}^{n}$ and $\mathcal{Y}\!=\!\{{y}_{i}\}_{i=1}^{n}$, respectively. The unified SFDA task is to learn a target model $f_{\theta_{t}}:\mathcal{X}\to\mathcal{Y}$ without any prior knowledge of the target domain not the relationships between the source and target domains (see Tab. I), given (i) the source model $f_{\theta_{s}}$ pre-trained over $\mathcal{X}_{s}$, $\mathcal{Y}_{s}$, and (ii) the unlabeled target data $\mathcal{X}$.
@@ 3.2 Causality in transfer learning with source domain
In transfer learning with source domain available, such as DG, and UDA, the causality relation hidden in the statistical dependence (between raw image $\boldsymbol{x}$ and available label $\boldsymbol{y}$) can be summarized to a Structural Causal Model (SCM). As depicted in Fig. 1(a), $\boldsymbol{x}$ is generated jointly by non-causal factors ${U}$ and causal factors ${S}$, while label $\boldsymbol{y}$ is determined by ${S}$ alone. Here, non-causal factors collectively represent the latent variables that generate the category-independent appearance of images, e.g., different domain styles and spurious dependence (Fig. 2); The causal factors present the ones determining the classification, e.g., the object shape.
@@
In summary, the key idea is when we change the probability distribution of non-causal factors (intervention) whilst keep the category (prediction consistency), the causality can be extracted. Often, domain related data augmentation is adopted as intervention, whilst the source labels ensure prediction consistency. However both conditions are unavailable in SFDA, making it inapplicable.
@@ 3.3 Latent Causality in Unified SFDA
To address the challenges outlined earlier, we propose an approach named Latent Causal Factors Discovery (LCFD), taking into account the following considerations. Instead of extracting causality in the raw input space $\boldsymbol{x}$, we shift our focus to the logit space ${\mathcal{A}}$: $\boldsymbol{a}=f_{\theta_{t}}(\boldsymbol{x})\in\mathcal{A}$. The choice of this space for causality analysis is crucial for two main reasons. First, at the presence of domain distribution gap during adaptation, using the logits essentially imposes a probability distribution perturbation, which could be considered as a special sort of intervention. This post-intervention property facilitates the extraction of causal factors [35, 30, 10]. Second, integration of additional information sources becomes simpler in this semantic space, without being constrained by distinct model architectures.
@@
Figure 3: Overview of our proposed LCFD framework: (a) Phase-1: Discovering the external causal factors $S_{e}$ in form of prompt context $\boldsymbol{v}_{e}$ from a frozen ViL model using our self-supervised information bottleneck algorithm; (b) Phase-2: Discovering the internal causal factors $S_{i}$ where the updated prompt context $\boldsymbol{v}_{e}$ is used to predict pseudo-labels as the prior information.
An overview of our LCFD is depicted in Fig. 1(b). It models non-causal $U$ and causal $S$ factors in a latent space ${\mathcal{A}}$. Uniquely, the causal part is further decomposed into internal $S_{i}$ and external $S_{e}$ components, consider that the information of source and target domain is not necessarily comprehensive. In contrast to conventional intervention driven learning, we focus on discovering $S$ from the intertwined $S$ and $U$ as
$\begin{split}P(\boldsymbol{y}|S)=P(\boldsymbol{y}|dis(S|\{U,S\})),~{}S=\{S_{i},S_{e}\}.\end{split}$ (2)
To make this process computationally tractable, we consider
@@
Independent Causal Mechanisms Principle [36]: The conditional distribution of each variable given its causes (i.e., its mechanism) does not inform or influence the other mechanisms.
This principle means $S_{i}$ and $S_{e}$ are independent in our context. So we can simplify our above formula as:
$\small\begin{split}P(\boldsymbol{y}|S)=P(\boldsymbol{y}|dis(S_{e}|\{U,S_{e}\}))\cdot P(\boldsymbol{y}|dis(S_{i}|\{U,S_{i}\})).\end{split}$ (3)
3.4 LCFD Design
We present a concrete design for realizing the LCFD as formulated above. An overview is depicted in Fig. 3. By Eq. (3), we implement the causality discovery in two parts: phase-1 for $S_{e}$ and phase-2 for $S_{i}$, as detailed below.

### v4

@@ Abstract
In the pursuit of transferring a source model to a target domain without access to the source training data, Source-Free Domain Adaptation (SFDA) has been extensively explored across various scenarios, including Closed-set, Open-set, Partial-set, and Generalized settings. Existing methods, focusing on specific scenarios, not only address a limited subset of challenges but also necessitate prior knowledge of the target domain, significantly limiting their practical utility and deployability. In light of these considerations, we introduce a more practical yet challenging problem, termed unified SFDA, which comprehensively incorporates all specific scenarios in a unified manner. In this paper, we propose a novel approach latent Causal factors discovery for unified SFDA (CausalDA). In contrast to previous alternatives that emphasize learning the statistical description of reality, we formulate CausalDA from a causality perspective. The objective is to uncover potential causality between latent variables and model decisions, enhancing the reliability and robustness of the learned model against domain shifts. To integrate extensive world knowledge, we leverage a pre-trained vision-language model such as CLIP. This aids in the formation and discovery of latent causal factors in the absence of supervision in the variation of distribution and semantics, coupled with a newly designed information bottleneck with theoretical guarantees. Extensive experiments demonstrate that CausalDA can achieve new state-of-the-art results in distinct SFDA settings, as well as source-free out-of-distribution generalization. Our code and data are available at https://github.com/tntek/CausalDA.
@@ 1 Introduction
Due to demands for privacy and information protection in public security and commercial competition, well-annotated training data are recognized as crucial for varying purposes, requiring stringent access control. However, in practical scenarios, the use of a source model (pre-trained on the source domain) without access to the actual source data, referred to as Source-Free Domain Adaptation (SFDA) [23], has become more feasible. This approach involves the challenges posed by strict data access controls and emphasizes the adaptation of a pre-trained model to new domains without relying on the availability of the original source data.
SFDA has been explored across various scenarios, as summarized in Tab. I. The most constrained scenario, termed Closed-set, assumes identical categories of interest in both the source and target domains, aiming to address covariate shift, such as domain-specific appearance distributions [29]. Generalized SFDA extends this by requiring the adapted model to perform well in both the source and target domains without forgetting the source domain [66]. Moving beyond the vanilla category identity constraint, Open-set SFDA [37] considers the target domain with additional categories, while Partial-set SFDA [4] takes the opposite direction. Both scenarios require handling the additional challenge of semantic shift.
@@
TABLE I: Summary of different SFDA settings. $C_{s}$$C_{t}$: The category set of the source/target domain.
@@
To address the aforementioned limitation, this study introduces a more realistic yet challenging problem, referred to as Unified Source-Free Domain Adaptation (Unified SFDA), with the goal of comprehensively addressing all the specific scenarios mentioned in a unified manner. To achieve this, we propose a novel approach latent Causal discovery of salient factors for unified SFDA (CausalDA), going beyond conventional statistical association learning of related variables by exploring the underlying causal relationships [38]. The causal mechanism discovered is expected to be more reliable in varying distributions and semantic contexts, providing a unified solution for different scenarios of SFDA.
In the absence of label supervision for both distributional and semantic variations, our CausalDA is specifically formulated using a structural causal model in the logits space. The latent causal factors are disentangled into two complementary parts: (i) external causal factors and (ii) internal causal factors, consider that the information of both domains is not necessarily complete. For the former, we leverage a pre-trained large Vision-Language (ViL) model with rich knowledge such as CLIP [42], which has been exposed to a vast amount of multimodality information sources. The latter is identified under the guidance of discovered external causal factors. The discovery of external and internal causal factors is alternated by designing a self-supervised information bottleneck with theoretical guarantees.
@@
    We propose a novel CausalDA approach for unified SFDA. Instead of learning the statistical description of problem reality as conventional methods do, CausalDA is formulated under a causality perspective. It attempts to reveal the potential causality between latent variables and model decisions, providing favorable robustness against both distributional and semantic shifts.
@@
    Extensive experiments demonstrate that CausalDA consistently achieves new state-of-the-art results on varying SFDA scenarios as well as source-free out-of-distribution generalization.
@@ 2.1 Source-free domain adaptation
Most of the prior SFDA methods consider the Closed-set setting where the source and target domains share the same classes, and the focus is on cross-domain distribution alignment. Existing methods exist in two categories. The first converts SFDA to Unsupervised Domain Adaptation (UDA) by constructing a pseudo-source domain [28], exploiting source prototype-guided data generation [55] or splitting source-like subset from target data [13]. The second follows the paradigm of self-supervised learning [31, 9], introducing self-guidance. Besides widely used pseudo-labels [31, 32], geometry information [64, 65, 54], contrastive data [68] and historical data [50] have also been exploited.
Recently, more practical settings have been studied. For example, the Generalized setting shifts aims to mitigate forgetting the source domain [66]. A representative approach is combining continual learning techniques with cross-domain adaptation, e.g., the domain attention-based gradient regulation [66] and source guidance constructed by historical information refining [50]. There exist works considering Open-set [33] and Partial-set [30] settings. For instance, to control the Open-set risk, a progressive balanced pseudo-labeling strategy [33], and Closed-set class prototypes are proposed [58]. Estimating the target data distribution helps with negative transfer reduction from the partial classes shift [26].
@@
Most recently, multimodal large models such as CLIP and ShareGPT [8] have been introduced to advance SFDA, substantially enhancing adaptation performance by leveraging generic knowledge in these models. Two strategies are adopted: (1) customizing and adapting this knowledge via mutual knowledge distillation [53, 67] or multi-teacher guidance [6] or multimodal space alignment [7]; and (2) further denoising ViL predictions [52]. Critically, we highlight a couple of key aspects that make CausalDA conceptually distinct from these SFDA methods. First, CausalDA is unique in design philosophy by learning a causal representation, in contrast to the statistical knowledge DIFO [53] and ProDe [52] both are designed to learn. As a consequence, the outreach and generalization can be further extended, leading to the first dedicated unified SFDA approach. Instead, DIFO and ProDe predominantly focus on the conventional closed-set setting. Further, this conceptual discrepancy leads to distinct leveraging of ViL knowledge. DIFO and ProDe both employ the generic ViL knowledge directly as supervision, fostering the establishment of statistical relationships, whilst CausalDA utilizes ViL knowledge to uncover the underlying causal factors for capturing the essence of visual recognition.
Existing work targets only specific SFDA settings, limiting its practical applicability and generality. This paper introduce a unified SFDA setting that incorporates all the pre-existing scenarios. With this new setup, we aim to foster more advanced SFDA methods that could generalize to a variety of problem settings without the need for designing and maintaining an array of distinct algorithms.
@@ 2.2 Causality methods in transfer learning
Causality methods aim for a robust function relation between random variables, which is insensitive to the extra variation/disturbing [45]. This property leads to recent popularity of this new paradigm in transfer learning, e.g., UDA [62], domain generalization (DG) [11], and out of distribution (OOD) [3]. These attempts aim to reduce non-causal factors by introducing artificial intervention. For example, the probability distribution change (used to represent the domain shift) and the semantic consistency constraint are jointly employed to implement this intervention. As source labels are available to ensure consistency, these methods emphasize constructing the distribution change in two lines. One explicitly generates augmented data to disturb the distribution, e.g., the non-linear augmentation and spatially-variable blending augmentation [36], inverse Fourier transformation augmentation [34], generative models-based cross-domain image style transformation [60], and causality-adjusted augmentation [72]. The other implicitly modifies the distribution by exploiting cross-domain data. For example, MatchDG [35] developed cross-domain contrastive learning where the input’s positive sample was randomly selected from samples with the same class. COR [61] adopted the variational inference encoder to infer unobserved causal factors from historical data, taking the time domain as a natural distribution change.
Compared with the existing works, the following three features distinguish CausalDA from them. (i) CausalDA does not rely on real labels that are indispensable for previous works. (ii) CausalDA discovers the causal factors instead of indirectly removing the non-causal ones. (iii) CausalDA builds the structural causal model at the logits level, whilst previous methods construct on raw data.
@@ 3.1 Unified SFDA problem
SFDA aims to transfer a model pre-trained on the source domain to a different but related target domain without labeling. Formally, let $\mathcal{X}_{s}\!=\!\{\boldsymbol{x}_{i}^{s}\}_{i=1}^{n_{s}}$ and $\mathcal{Y}_{s}\!=\!\{{y}_{i}^{s}\}_{i=1}^{n_{s}}$ be the source samples and their truth labels. Similarly, the unlabeled samples and the truth target labels are $\mathcal{X}\!=\!\{\boldsymbol{x}_{i}\}_{i=1}^{n}$ and $\mathcal{Y}\!=\!\{{y}_{i}\}_{i=1}^{n}$, respectively. The unified SFDA task is to learn a target model $f_{\theta_{t}}:\mathcal{X}\to\mathcal{Y}$ without any prior knowledge of the target domain not the relationships between the source and target domains (see Tab. I), given (i) the source model $f_{\theta_{s}}$ pre-trained over $\mathcal{X}_{s}$, $\mathcal{Y}_{s}$, and (ii) the unlabeled target data $\mathcal{X}$.
@@ 3.2 Causality in transfer learning with source domain
In transfer learning with available source domain, such as DG and UDA, the causality relation hidden in the statistical dependence (between raw image $\boldsymbol{x}$ and available label $\boldsymbol{y}$) can be summarized to a Structural Causal Model (SCM). As depicted in Fig. 1(a), $\boldsymbol{x}$ is generated jointly by non-causal factors U{U} and causal factors S{S}, while label $\boldsymbol{y}$ is determined by S{S} alone. Here, non-causal factors collectively represent the latent variables that generate the category-independent appearance of images, e.g., different domain styles and spurious dependence (Fig. 2); The causal factors present the ones determining the classification, e.g., the object shape.
@@
In summary, the key idea is that when we change the probability distribution of non-causal factors (intervention) whilst keeping the category (prediction consistency), the causality can be extracted. Often, domain-related data augmentation is adopted as an intervention, whilst the source labels ensure prediction consistency. However, both conditions are unavailable in SFDA, making it inapplicable.
@@ 3.3 Latent causality in unified SFDA
To address the challenges outlined earlier, we propose an approach named latent Causal factors discovery (CausalDA), taking into account the following considerations. Instead of extracting causality in the raw input space $\boldsymbol{x}$, we shift our focus to the logit space $\bm{A}$}: $\boldsymbol{a}=f_{\theta_{t}}(\boldsymbol{x})\in\bm{A}$. The choice of this space for causality analysis is crucial for two main reasons: First and foremost, from a causal perspective, using a post-intervened space enhances causality capture by incorporating a probability distribution shift into the data [35, 61, 38]. In our case, the logit space is post-intervened due to the inherent distribution shift between the in-training target model (initialized with the pre-trained source domain model) and the target domain. Specifically, the target model’s logits for the target domain data reflect this domain shift, making them a particular form of post-intervention in this context.
Furthermore, the logit space is both highly semantic and compact, making it more efficient to manipulate. Since logits encode rich semantic information about class relationships, operating in this space allows for better generalization and transferability across domains. Additionally, its compact representation reduces computational complexity, facilitating more efficient learning and adaptation. These properties collectively make the logit space a more effective choice for our approach. Tab. II presents an analogy of CausalDA and the existing causal learning approach.
@@
Approach Intervention Consistency
Existing [10, 11, 61] Data augmentation Supervised
CausalDA Using logit Unsupervised
TABLE II: Analogy of causality learning.
Figure 3: Overview of our CausalDA framework: (a) Phase-1: Discovering the external causal factors $S_{e}$ in form of prompt context $\boldsymbol{v}_{e}$ from a frozen ViL model using our self-supervised information bottleneck algorithm; (b) Phase-2: Discovering the internal causal factors $S_{i}$ where the updated prompt context $\boldsymbol{v}_{e}$ is used to predict pseudo-labels as the prior information.
An overview of our proposed CausalDA framework is illustrated in Fig. 1(b). It models both non-causal factors $U$ and causal factors $S$ within a latent space A{A}. Notably, the causal component $S$ is further decomposed into internal ($S_{i}$) and external ($S_{e}$) sub-components to facilitate a more complete and interpretable modeling of causality.
This internal-external causality decomposition is motivated by the recognition that identifying all relevant causal relationships solely from the source and target domains is often infeasible due to limited training data and domain-specific knowledge. To address this, we introduce a pretrained ViL model, such as CLIP, as an external knowledge source to aid in the discovery of causal relationships. This design leads to two types of causality: (i) Internal causality referring to causal factors that can be inferred from the source domain model and the target domain data; (ii) External causality, additional complements captured from the ViL model with broad, many-domain information. Such a collaborative design enables a more holistic understanding of causal structures, as empirically validated in Sec.4.2 and Sec.4.7.2.
In contrast to conventional intervention-driven learning, we focus on discovering $S$ from the intertwined $S$ and $U$ as
$\begin{split}P(\boldsymbol{y}|S)=P(\boldsymbol{y}|dis(S|\{U,S\})),~S=\{S_{i},S_{e}\},\end{split}$ (2)
where $dis(S|\{U,S\}))$ means conducting discover operation on $\{U,S\}$ to decompose $S$. To make this process computationally tractable, we consider
@@
Independent Causal Mechanisms (ICM) Principle [41]: The conditional distribution of each variable given its causes (i.e., its mechanism) does not inform or influence the other mechanisms.
As in Sec.3.2, the lack of access to the source domain prevents the use of conventional causal learning methods. Inspired by Principle 1, we assume conditional independence between external and internal causal factors to enable a divide-and-conquer learning approach. Under this perspective, we propose the following factorized approximation as:
$\small\begin{split}P(\boldsymbol{y}|S)\approx P(\boldsymbol{y}|dis(S_{e}|\{U,S_{e}\}))\cdot P(\boldsymbol{y}|dis(S_{i}|\{U,S_{i}\})).\end{split}$ (3)
3.4 CausalDA design
We present a concrete design to carry out CausalDA, as depicted in Fig. 3. By Eq. (3), we implement the causality discovery in two successive phases: phase-1 for $S_{e}$ and phase-2 for $S_{i}$, as detailed below.
Figure 4: Realizing our self-supervised information bottleneck w.r.t. Theorem 9. $\rm{T}_{1}$ is estimated by Variational Mutual Information (VMI) with Gaussian distribution assumption $\mathcal{N}(Z,\Sigma_{Z})$. $\rm{T}_{2}$ is computed by Probabilistic Mutual Information (PMI) where function $f_{w}$ integrates min-max optimization, using covariance matrix $\Sigma_{Z}$ from $\rm{T}_{1}$ as weighting parameters $w$.